### v1

# Time-Optimized Safe Navigation in Unstructured Environments through Learning Based Depth Completion

Abstract
Quadrotors hold significant promise for several applications such as agriculture, search-and-rescue, and infrastructure inspection. Achieving autonomous operation requires systems to navigate safely through complex and unfamiliar environments. This level of autonomy is particularly challenging due to the complexity of such environments and the need for real-time decision-making—especially for platforms constrained by size, weight, and power (SWaP), which limits flight time and precludes the use of bulky sensors like Light Detection and Ranging (LiDAR) for mapping. Furthermore, computing globally optimal, collision-free paths and translating them into time-optimized, safe trajectories in real time adds significant computational complexity. To address these challenges, we present a fully onboard, real-time navigation system that relies solely on lightweight onboard sensors. Our system constructs a dense 3D map of the environment using a novel visual depth estimation approach that fuses stereo and monocular learning-based depth, yielding longer-range, denser, and less noisy depth maps than conventional stereo methods. Building on this map, we introduce a novel planning and trajectory generation framework capable of rapidly computing time-optimal global trajectories. As the map is incrementally updated with new depth information, our system continuously refines the trajectory to maintain safety and optimality. Both our planner and trajectory generator outperforms state-of-the-art methods in terms of computational efficiency and guarantee obstacle-free trajectories. We validate our system through robust autonomous flight experiments in diverse indoor and outdoor environments, demonstrating its effectiveness for safe navigation in previously unknown settings.
{IEEEkeywords}
Aerial Systems: Perception and Autonomy, Collision Avoidance, RGB-D Perception, Vision-Based Navigation
Video: https://www.youtube.com/watch?v=vC5Dw1idMug
1 Introduction
Method Devo[1] MINCO [2] Saviolo [3] OUR Rappids [4] Lee[5] Super [6] Zhou [7]
Sensors 1x RGB 1x RGBD 1x RGBD 1x RGBD 2x Stereo Pairs 2x RGBD 1x LiDAR 1x LiDAR
Total Sensor Mass 5 g 15 g 72 g 100 g 127 g 200 g 265 g 840 g
Sensor Range N/A 3 m 8 m 10 m 3 m 6 m 40 m 100 m
Safety Guarantee No No Yes Yes Yes Yes Yes Yes
Traj. Planning Learning Global Local Global Local Local Global Local
Table 1: Autonomous Navigation Systems Comparisons. Safety Guarantees in our use case is defined as the trajectory solved can never intersect with an obstacle or closer than some safety distance. It should be noted that Super [6] trajectory generation is based on MINCO.
Figure 1: A map of our environment to the left along with the flight path. Right is the color image and our generated camera depth
[…]
Autonomous Micro Aerial Vehicles (MAVs) offer valuable solutions across a range of application scenarios such as agriculture [8], search-and-rescue [9], and infrastructure inspection [10]. To perform these maneuvers, the aerial vehicles have to locate the obstacles in real time while concurrently generating and executing safe and obstacle free trajectories.
Many existing systems rely on heavyweight sensors such as Light Detection and Ranging (LiDAR) to ensure collision-free navigation [6]. However, these sensors are often incompatible with the Size, Weight, and Power (SWaP) constraints of small-scale aerial vehicles, limiting flight time and agility. While lightweight alternatives like cameras are more suitable for SWaP-constrained platforms, they frequently struggle with noisy and inaccurate obstacle localization. Moreover, computing globally optimal, obstacle-free paths in large-scale environments is computationally intensive, and transforming these paths into time-optimal, safe trajectories in real time introduces additional complexity.
To address these challenges, we propose an autonomous quadrotor navigation framework that enables real-time exploration and mapping of unknown indoor and outdoor environments. The system operates using a lightweight sensor suite consisting of cameras, an IMU, and GPS (when available), making it well-suited for SWaP-constrained platforms. Our navigation stack can generate high quality depth, convert it to an accurate map, and autonomously plan safe trajectories around obstacles illustrated in Fig. 1. In this context, safety refers to the system’s ability to provide guarantees for the planned trajectories by ensuring they avoid collisions with obstacles and stay within operational limits, based on the robot’s current environment map.
Contributions. The contributions of this work are
  • We develop a visual perception algorithm capable of generating dense and low-noise point clouds. Our method combines monocular, learning-based depth estimation with a traditional commercial off-the-shelf stereo matching solution. As a result, it produces depth estimates †with longer range, lower noise, and improved structural consistency compared to standard commercial solutions.
  • We propose the first real-time implementation in unknown environments for shortest path on the graph of convex sets (our path planning method). Our path planning method computation scales well with the size and complexity of maps outperforming other popular methods like Rapidly-exploring Random Tree (RRT).
  • We develop a unique time-optimized trajectory generation method with safety guarantees. We compare our trajectory generation method to MINCO [11] another time-optimized trajectory method used in multiple popular works [6, 2, 11] . Our method generates global trajectories more quickly while maintaining comparable optimality and ensuring safety.
  • We extensively validate our navigation system in both indoor (GPS denied) and outdoor environments for exploration tasks under varying lighting conditions and harsh winds of up to $15$\mathrm{m}\mathrm{p}\mathrm{h}$$, demonstrating reliable performance over areas of up to $60$\mathrm{m}$\times 40$\mathrm{m}$\times 3$\mathrm{m}$$.
2 Related Works
Autonomously Navigating environment requires two major components. First the vehicle needs to be able to localize the obstacles in the environments. Next, the vehicle needs to intelligently plan a path that avoids obstacles. Several related works featuring both components are listed in Table 1. Overall, our approach offers the best sensing range among vision-based methods using lightweight cameras, while maintaining safety guarantees for the planned trajectories. In the following, we first focus on describing relevant works in the mapping area and then planning approaches.
Mapping. There are two predominant sensors that are currently employed to generate environment depth onboard a quadrotor: cameras or LiDAR. Although other sensors exist, such as radar[12] or ultrasounds [13], they are far less prevalent on drones due to their size or low resolution. Most works using LiDARs [14, 7, 6, 15] are able to achieve strong mapping and localization performance. However, the LiDAR’s weight ranges from $250\leavevmode\nobreak\ $\mathrm{g}$$ to $1000\leavevmode\nobreak\ $\mathrm{g}$$ which limits their use on smaller aerial platforms. In contrast, stereo cameras offer a much lighter payload, typically less than $100\leavevmode\nobreak\ $\mathrm{g}$$. However small stereo camera pairs have low maximum depth sensing range, typically $3\leavevmode\nobreak\ $\mathrm{m}$-6\leavevmode\nobreak\ $\mathrm{m}$$ [2, 4, 5]. Learning based monocular depth estimation methods [16] offer an exciting new paradigm. Monocular depth estimation inferences depth based on the image’s structure by learning the mapping from images to depth on large datasets. Unfortunately, while the depth structure is relatively consistent, the scale of the depth is wildly off for objects outside the training set. One popular paradigm for addressing this problem is called depth completion, where a sparse metric depth supplements monocular information to create a dense structurally consistent metric depth. A supplemental sensor like a radar [17] can be fed into a larger neural network to infer the scale. However, this solution involves large datasets of unified sensor data to create this depth completion method therefore limiting their deployment. Unlike those methods, in the proposed work we develop an optimization-based scaling method using our sparse depth generated from our RGB-D camera similar to [3], but we employ inverse depth obtaining more accurate and consistent results. This method does not require retraining and is applicable to a variety of settings giving us great flexibility for deployment and switching the components in this system.
Figure 2: System block diagram. The perception and planning/control pipelines are run onboard the Orin NX. Our autopilot is the Pixracer Pro running Px4 firmware. Stars represent the start and goal for the trajectory planning. The start information comes from the robot state, and the goal is user given or done autonomously with the exploration algorithm.
Planning. There are two main planning approaches for obstacle avoidance: local planning and global planning. Local planning [3, 4, 5, 18] is favored for its fast computation, making it well-suited for high-speed reactive avoidance. However, since it only considers local information, it struggles to navigate dead ends or non-convex obstacles. To handle such situations, global information is typically integrated at a higher level of abstraction through a separate global planner. Global planning involves computing a collision-free path across the entire environment map and refining it into a feasible trajectory [2, 19, 20, 21, 22]. Global path generation methods generally fall into two categories: sampling-based and search-based approaches. These methods are also applicable to local planning if we restrict the map size to a local region. However, in the following, we will evaluate the performance of the following method within the context of a large-scale global map. Sampling-based methods, such as Rapidly-exploring Random Trees (RRT) [22] and the Probabilistic Roadmap Method (PRM) [21], generate paths by randomly sampling the space and connecting traversable states to form a space filling tree. These methods scale well to high-dimensional spaces and avoid space discretization. However, they lack optimality guarantees and have unbounded computation times. In contrast, search-based methods, including A* [23, 24] and jump point search [25], operate on a discretized representation of the space, often uniform voxels. These methods provide optimal solutions, have bounded solving times, and can determine when a goal is unreachable. However, their scalability is poor in large environments as a result of the cubic growth of the search space. To address the poor space scalability, other works abstract the occupancy as a graph of convex sets [26, 27, 28]. The graph of convex sets representation converts the free space to a sparse set of convex objects, axis-aligned cuboids for our implementation. Since each element in the graph represents a full volume of uniform voxels, the number of elements in the graph is greatly decreased which simplifies the search problem. However, previous works assume a fully known environment and use preprocessing to generate the graph. Our method is the first to apply this technique in real-time with a dynamically updating map. Solving the path planning problem on this graph is the shortest path problem on a graph of convex sets [26, 27, 28]. Solving the path is NP-Hard however there exists a few approximation methods using either mixed integer convex programming (MICP) [26] or coverts the convex sets to singular vertices [28] and use a shortest path on a singular graph like Djikstra, or A*. Mixed integer programming is close to optimal solution but slow to compute. While graph based search methods are quicker to compute, but converting a set of convex sets to set of vertices is another optimization problem. [28] assumes that the map is known and solves this conversion in pre-processing. Since our application involves real-time flight, long pre-processing steps are not feasible. To address this, our approach assumes that the all the edges of our graph of convex set are equal and constant and uses the minimum distance between each convex set and the goal as a heuristic. This design significantly improves the scalability of our path planning method in terms of computation time.
Once a set of safe flight corridors (SFC) is determined, computing a feasible trajectory becomes a well-studied problem [29, 19, 24]. One of the most popular methods is MINCO[11, 2] which uses an unconstrained optimization with polynomial splines to plan through SFC. MINCO has shown better optimality in terms of time and control effort than other baselines [30, 31, 32] in MINCO’s own experimental evaluation. Additionally, it has show quick computation times in many of its application works [33, 6, 2]. We propose a constrained optimization guaranteeing our constraints are respected unlike [11] . Second, we replace its polynomial spline basis with Bernstein polynomial splines [34]. This change of basis allows continuous safety enforcement without discretization [2], or additional post-processing for collision checks and replanning in proposed other works [22, 21]. Additionally unlike other works using Bernstein polynomial splines [34, 35], our work does not directly optimize control points or coefficients but the start and end points of each spline reducing the complexity of the optimization. Our results show that our method takes less computation time [11] while maintaining comparable optimality and guaranteeing safety.
3 Methodology
In this section, we describe our navigation approach made up of two major parts as seen in Fig. 2, a mapping and planing/control pipeline. First, we require our robust mapping pipeline which generates both an occupancy map and graph of free space. This graph of free space is then fed into our robust planning and control algorithm design.
3.1 Mapping
In this section, we describe our depth generation method. We utilize an RGB-D camera that outputs temporally and spatially synchronized color and depth images. Our goal is to take the noisy depth image created by stereo and combine it with color information to create a less noisy and structurally consistent depth. First, we feed our color image ($W\times H\times 3$ dimensions) is fed to a monocular depth estimation network, DepthAnythingv2 [16], which generates an unscaled inverted depth $\mathbf{d}_{m}$ with units $\mathrm{m}\mathrm{m}^{-1}$. Monocular depth is an ill-posed problem. The structure created by the depth is relatively accurate which still makes it useful. However, the scale information is highly inaccurate. Conversely, stereo depth, $\mathbf{d}_{s}$ in $\mathrm{m}\mathrm{m}$, has a reliable scale but lacks structure and is filled with speckle noise and holes. Ideally, we want to fuse the monocular structure from the monocular depth with the scale from the stereo. To achieve this objective, we rescale our monocular depth with the following optimization problem
$\begin{split}\min_{\alpha_{2,1,0}}\sum_{j=0}^{W}\sum_{i=0}^{H}&\left(\frac{1}{\mathbf{d}_{s}[i,j]}-\alpha_{2}\mathbf{d}_{m}[i,j]^{2}-\alpha_{1}\mathbf{d}_{m}[i,j]-\alpha_{0}\right)^{2}\\ \mathbf{d}_{c}[i,j]&=\frac{1}{\alpha_{2}\mathbf{d}_{m}[i,j]^{2}+\alpha_{1}\mathbf{d}_{m}[i,j]+\alpha_{0}},\end{split}$ (1)
where $\mathbf{d}_{c}$ is the completed depth image in $\mathrm{m}\mathrm{m}$, $[\alpha_{2},\alpha_{1},\alpha_{0}]$ are scaling coefficients, and $i$ and $j$ are the index terms for the depth images. We choose a square approximation because it has been shown to be optimal in literature [3] for this problem. The scaling coefficients are solved using the least squares regression algorithm [36]. We perform our optimization on the $\mathrm{m}\mathrm{m}$^{-1}$ scale, as it consistently yields superior results compared to the $\mathrm{m}\mathrm{m}$ scale. In least squares regression, the largest values disproportionately influence the optimization outcome. When using the $\mathrm{m}\mathrm{m}$^{-1}$ scale, points closer to the camera have a greater impact on the optimization. Closer points to the camera benefit from higher stereo depth accuracy than distant points. In contrast, optimization on the $\mathrm{m}\mathrm{m}$ scale is dominated by distant points. As a result, the $\mathrm{m}\mathrm{m}$^{-1}$ scale is more suitable for achieving accurate depth optimization. The completed depth $\mathbf{d}_{c}$ and vehicle’s pose is fed to a mapping algorithm, Nvblox [37], to generate the occupancy map. After the occupancy is generated, we extract the free space and perform the cuboid decomposition described in the following.
3.2 Cuboid Decomposition and Path Selection
To generate an obstacle-free, dynamically feasible trajectory, our approach start by generating a set of Safe Flight Corridors (SFC) connecting the start and goal. Most existing methods [11, 19, 20] first compute a obstacle free path and then create SFC around the path. Our method instead reverses this procedure converting the occupancy map to a graph of SFC, and then finding the subset of SFC that links start and goal. We call our clustering approach to turn the occupancy map into a graph of safe flight corridors through cuboid decomposition. The advantage of our approach is that the clustering can be performed at the map generation step in parallel to trajectory planning before a goal is given. Additionally, once the clustering is complete, solving the subset of safe flight corridors is faster scaling with the number of SFC in the graph rather than size of the map.
Algorithm 1 Cuboid Decomposition
0:   Occupancy Grid
0:   A graph of cuboids with a set of vertices $\mathbb{V}$ that represent free space sets and edges $\mathbb{E}$ that represent their connectivity
  Initialization: Mark all free space as uncovered.
  while There exists an uncovered node do
     Node Selection: Select an uncovered and unoccupied node, as shown in the search step (Fig. 3a).
     Cover: Expand a cuboid $v_{k}$ from the selected node in all directions until it touches obstacles on each side seen in Fig. 3b.
     Add Vertices $\mathbb{V}=v_{k}\cup\mathbb{V}$
     Mark: Mark all voxels within the added cuboid as covered.
  end while
  Solve Edges
  for $v_{i}\in[\mathbb{V}]$ do
     for $v_{j}\in[\mathbb{V}]$ do
        if $v_{i}\cap v_{j}\neq\varnothing$ then
           Make Edge $e_{k}=[v_{i},v_{j}]$
           Add Edge $\mathbb{E}=e_{k}\cup\mathbb{E}$
        end if
     end for
  end for
  return  Graph of Convex Sets $\mathbb{V}$ and $\mathbb{E}$
Figure 3: Cuboid Decomposition: First discover an unexplored and unoccupied node. Next create a cuboid to cover the free space expanding in all directions.
In the literature, cuboid decomposition can be viewed as a covering problem, where the goal is to cover the free space with boxes, as described in [38]. Our proposed cuboid decomposition algorithm is illustrated in Fig. 3. The procedure involves the following steps described in our Algorithm  1. In general, we first repeatedly create cuboids to cover the free space. Subsequently, we construct the edges of our graph by checking if each pair is connected. Each vertex of this graph represents a cuboid of free space and each edge represents an adjacency. This graph-based representation is updated repeatedly, independently of path planning. To extend this decomposition into $2.5$D, we divide the space into several $2$D layers and repeat the process for each layer.
Once we have a graph representation of the free space, we solve for our path. Rather than an explicit path, we are selecting a set of SFC that connect our start and goal positions. We implement a modified A* algorithm to achieve our goal. The edge weight of each corridor is $1$. The lower bound heuristic is the minimum distance between a point in the cuboid node and the goal, assuming that no obstacles exist. As the number of cuboids is significantly lower than the nodes in an occupancy grid, A* can be solved in less than $1\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ while providing a near-optimal set of safe flight corridors as seen in Fig. 4c. Solving for a true optimal set of safe flight corridors is a NP-Hard problem [28]. Conversely, this algorithm’s objective is to minimize the number of SFC which in turn corresponds to minimizing the number of optimization variables for our trajectory generation.
3.3 Bernstein Time-Optimized Trajectory Planning
Figure 4: Trajectory Planning Pipeline. Cuboid Decomposition refers to the procedure detailed in Fig. 3 and Algorithm 1. Safe Flight Selection refers to finding the minimum number of safe flight corridors that connect a start and a goal. The light gray area is where the cuboids are covering the free space. Darker gray areas are just where the safe flight corridors intersect. Red dots are the optimization variable which are initialized in step (d) and finally optimized in step (e).
We solve for a trajectory that connects a start condition of position, velocity, and acceleration $[\mathbf{x}_{0}^{\top},\mathbf{v}_{0}^{\top},\mathbf{a}_{0}^{\top}]^{\top}$ with a goal position $\mathbf{x}_{f}$ given a set of safe flight corridors. We introduce the formalism for a single-dimensional case. To extend this to multiple dimensions, it is sufficient to repeat the constraints for each dimension. We formulate our trajectory using Bernstein polynomial splines. Each spline is assigned to a single safe flight corridor. First, we define each spline as follows
$\begin{split}B(t)&=\begin{cases}\text{$B_{0}\left(t\right)$}&\quad\text{if $t\in\left[0,t_{0}\right]$}\\ \text{$B_{1}\left(t-t_{0}\right)$}&\quad\text{if $t\in\left[t_{0},t_{0}+t_{1}\right]$}\\ \hskip 10.0pt\vdots\\ \text{$B_{f}\left(t-\sum\limits_{i=1}^{f-1}t_{i}\right)$}&\quad\text{if $t\in\left[\sum\limits_{i=1}^{f-1}t_{i},\sum\limits_{i=1}^{f}t_{i}\right]$}\\ \end{cases},\\ \end{split}$ (2)
where $t_{i}$ represents the time spent in the safe flight corridor $i$. $f$ is the number of safe flight corridors. Each spline $B_{i}(t)$ is defined as the following function
$\begin{split}B_{i}(t)&=\sum_{k=0}^{5}c^{k}_{i}b^{5}_{k}\left(\frac{t}{t_{i}}\right),\\ \end{split}$ (3)
where $c^{k}_{i}$ is the coefficient of the polynomial, $k$ represents the coefficient number and $i$ is the spline number. In literature, an alternative name for the coefficient is control points when describing Bernstein polynomials. $b^{5}_{k}(t)$ represents the basis function defined in [34] for order $5$ and 1-D polynomial.
However instead of structuring the problem for coefficients, we structure our optimization as solving for a set of waypoints $\mathbf{w}=[\mathbf{w}_{0},\mathbf{w}_{1},...,\mathbf{w}_{f}]^{\top}$ and traversal times $\mathbf{t}=[t_{0},t_{1},...t_{f}]^{\top}$ We define a waypoints as a $\mathbf{w}_{i}\in\mathbb{R}^{3\times 1}$ with elements $\mathbf{w}_{i}=\begin{bmatrix}B_{i}(0)&B_{i}^{\prime}(0)&B_{i}^{\prime\prime}(0)\end{bmatrix}^{\top}$ where $B^{\prime}=\frac{dB(t)}{dt}$ is the velocity, and $B^{\prime\prime}=\frac{d^{2}B(t)}{dt^{2}}$ is acceleration. Each spline is completely defined by the start and end waypoint on the curve. This relationship between coefficients and waypoints is defined as
$\begin{bmatrix}c^{0}_{i}\\ c^{1}_{i}\\ \vdots\\ c^{4}_{i}\\ c^{5}_{i}\\ \end{bmatrix}=\begin{bmatrix}1&0&0&0&0&0\\ 1&\frac{1}{5}t_{i}&0&0&0&0\\ 1&\frac{2}{5}t_{i}&\frac{1}{20}(t_{i})^{2}&0&0&0\\ 0&0&0&1&-\frac{2}{5}t_{i}&\frac{1}{20}(t_{i})^{2}\\ 0&0&0&1&-\frac{1}{5}t_{i}&0\\ 0&0&0&1&0&0\\ \end{bmatrix}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix},$ (4)
Building on this structure, we formulate the trajectory generation problem with the following optimization
$\min_{\mathbf{w}_{[0,...,f]},\mathbf{t}_{[0,...,f]}}$ $\leavevmode\nobreak\ \sum^{f}_{i=0}t_{i},$ (5)
[…]
$\mathbf{w}_{f}$ $=\begin{bmatrix}\mathbf{x}_{f}^{\top}&0&0\end{bmatrix}^{\top},$
[…]
The matrices $\mathbf{y},\mathbf{G},\mathbf{z}$ are safety constraints based on eq. (6). $\begin{bmatrix}\mathbf{x}_{0}^{\top}&\mathbf{v}_{0}^{\top}&\mathbf{a}_{0}^{\top}\end{bmatrix}^{\top}$ is the quadrotor’s current position, velocity, and acceleration respectively, $\mathbf{x}_{f}$ is the goal position. This results in a time-optimized trajectory in Fig. 4e.
Structuring the problem using waypoints instead of coefficients improves our optimization in three ways. First, optimizing on waypoints reduces the number of optimization variables compared to coefficients when increasing the number of splines. For example, with $f$ number of splines of order $5$, the number of optimization variables using waypoints is $3(f+1)$ while using coefficients it requires $6f$. Next, the gradients are much better conditioned on functions like $t_{i}^{2}$ and $t_{i}$ as opposed to optimizing over coefficients which involve the inverse, $\frac{1}{t_{i}}$ or $\frac{1}{{t_{i}}^{2}}$. Finally, it is easier to initialize our optimizer because waypoints have a better physical interpretation with respect to coefficients.
Subsequently, we define our safety constraints. The safe flight corridors are defined by their position constraints lower $\mathbf{l}=[l_{0},l_{1},...l_{f}]^{\top}$ and upper $\mathbf{u}=[u_{0},u_{1},...u_{f}]^{\top}$ for each corridor. With these bounds, we formulate our safety constraints from eq. (5), $\mathbf{y}$, $\mathbf{G},\mathbf{z}$, based on the convex hull property of Bernstein polynomials [34]. The convex hull property states bounds on the coefficients are equivalent to bounds on the full trajectory defined formally as
$\begin{split}l_{i}\leq\mathbf{A}_{i}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix}\leq u_{i}&\implies l_{i}\leq B_{i}(t)\leq u_{i}\hskip 5.0pt,\\ -\mathbf{v}_{m}\leq\mathbf{D}^{(1)}_{i}\mathbf{A}_{i}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix}\leq\mathbf{v}_{m}&\implies-\mathbf{v}_{m}\leq B^{\prime}_{i}(t)\leq\mathbf{v}_{m},\\ \underbrace{-\mathbf{a}_{m}\leq\mathbf{D}_{i}^{(2!)}\mathbf{A}_{i}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix}\leq\mathbf{a}_{m}}_{\text{Optimization Constraint}}&\implies\underbrace{-\mathbf{a}_{m}\leq B^{\prime\prime}_{i}(t)\leq\mathbf{a}_{m}}_{\text{Physical Result}},\\ \end{split}$ (6)
where $\mathbf{v}_{m}$, $\mathbf{a}_{m}$ is the maximum velocity and acceleration respectively. $\mathbf{A}_{i}$ is the matrix to map waypoints to coefficient defined in eq. (4). $\mathbf{D}_{i}^{(m!)}\in\mathbb{R}^{(6-m)\times 6}$ is a differentiation matrix that converts the coefficient to the $m$-th time derivative’s coefficient. The matrix $\mathbf{D}_{i}^{(m!)}$ is composed from $\mathbf{D}_{i}^{(m)}\in\mathbb{R}^{(6-m)\times(7-m)}$ as
$\begin{split}\mathbf{D}^{(m)}_{i}&=\frac{6-m}{t_{i}}\begin{bmatrix}-1&1&0&0&\ldots\\ 0&-1&1&0&\ldots\\ &&\ddots&\ddots&\\ \end{bmatrix},\\ \mathbf{D}^{(m!)}_{i}&=\mathbf{D}^{(1)}_{i}\mathbf{D}^{(2)}_{i}\ldots\displaystyle\mathbf{D}^{(m)}_{i}.\end{split}$ (7)
We then generate the initial trajectory $\mathbf{w}_{\text{init}}$. This trajectory is a straight line trajectory between the start point and the center of the intersections of each safe flight corridors visualized in Fig. 4d or mathematically defined as
$\begin{split}\mathbf{x}_{i}&=\frac{1}{2}(\max[l_{i},l_{i+1}]+\min[u_{i},u_{i+1}]),\\ \mathbf{w}_{i}&=\begin{bmatrix}\mathbf{x}_{i}^{\top}&0&0\\ \end{bmatrix}^{\top},\hskip 10.0pt1\leq i\leq f-1,\\ \mathbf{w}_{\text{init}}&=\begin{bmatrix}\mathbf{w}_{0}^{\top}&\mathbf{w}_{1}^{\top}&\ldots&\mathbf{w}_{i}^{\top}&\ldots&\mathbf{w}_{f}^{\top}\\ \end{bmatrix}^{\top},\\ \mathbf{t}_{\text{init}}&=\begin{bmatrix}\frac{||\mathbf{x}_{1}-\mathbf{x_{0}}||}{||\mathbf{v}_{m}||}&\frac{||\mathbf{x}_{2}-\mathbf{x_{1}}||}{||\mathbf{v}_{m}||}&\ldots&\end{bmatrix}^{\top}.\end{split}$ (8)
After we acquire our initialization, we complete the optimization described in eq. (5) to solve for the the time-optimized trajectory. Since our camera has a limited FoV and is forward facing, we need to constantly align our camera with the axis of motion. We implement a simple planner for determining the $\psi_{d}$ or desired yaw of the vehicle.
$\psi_{d}(t)=\begin{cases}\psi_{c}&\quad\text{if $||v_{y}||,||v_{x}||<1e-1$}\\ \psi_{c}-\psi_{m}&\quad\text{if $\arctan(v_{y},v_{x})-\psi_{c}<-\psi_{m}$}\\ \psi_{c}+\psi_{m}&\quad\text{if $\arctan(v_{y},v_{x})-\psi_{c}>\psi_{m}$}\\ \arctan(v_{y},v_{x})&\quad\text{else}\end{cases}$ (9)
where $\psi_{c}$ is the current yaw angle. $\psi_{m}$ is a limit on the maximum turning speed. $v_{y}$ and $v_{x}$ are the quadrotors current velocity projected to the world’s $x$ and $y$ axes.
To handle newly observed obstacles, we implement a repeated replanning procedure. First, we estimate if our currently planned trajectory intersects with the updated map’s obstacles within the next $10\leavevmode\nobreak\ $\mathrm{s}$$. If no collision is detected, then we continue with this flight plan. Otherwise, we plan a new trajectory with the updated map to our goal. In the case where no new trajectory can be solved, we stop the quadrotor (for example the goal is in a wall). This check of deciding whether to continue, stop, or replan is repeated at a rate of $20\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$.
Figure 5: The reference and estimate path of the $80\%$ exploration run including the tracking error.
Figure 6: Outdoor Exploration: We show our exploration paths given different completion goals in the environment.
Figure 7: The generated map by the quadrotor after an exploration run.
Figure 8: Depth Comparison. D455 Depth in blue is the stereo depth generated directly from the camera. Our Depth is the depth completion method described in our methodology.
3.4 Exploration
With our navigation pipeline, exploration is framed as goal selection. As exploration is not our primary contribution and is instead an application for navigation, we employ the Next Best View (NBV) algorithm [39] for its simplicity. First, the user defines a space to explore the environment. Within this space, we maintain a discretized voxel map representing whether this position has been viewed. Initially, the entire map is marked as unexplored. As the camera observes the environment, visible voxels are updated to be explored, gradually constructing an understanding of the area. To start the exploration, the robot performs a $360^{\circ}$ scan at the beginning of each mission. This maneuver provides initial information for our exploration and obstacle avoidance. Goal selection is handled using a sampling-based approach called NBV. We randomly sample $300$ candidate viewpoints within the voxel grid and evaluate how many unexplored voxels each viewpoint would newly reveal. We then select the viewpoint that maximizes this expected information gain. Navigation proceeds toward the selected goal until the goal is reached or found unreachable. Afterward, the quadrotor stops and selects another goal to continue the exploration. This exploration loop continues until an user discovery threshold is reached.
4 Experimental Results
For our experiments, we want to show our navigation framework is effective in both indoor and outdoor scenarios, and that the individual components compare well to standard baselines. First, we stress our system stack by continuously exploring an outdoor arena to different completion levels. Next, we compare the trajectory generation to MINCO[11], a popular baseline. Our method achieves greater computational efficiency while producing trajectories that are comparable in quality to those generated by this approach. Third, we show our depth generation method is superior to the depth output from the stereo baseline. These improvements include less noise, better structure, and a deeper range. Finally, our approach is capable of generalizes to indoor GPS-denied environments by trying more scenarios inside and challenging navigation tasks where the quadrotor has to escape a dead-end
4.1 System Setup
For our experiments, we consider a $1.3\leavevmode\nobreak\ $\mathrm{k}\mathrm{g}$$ drone with a thrust to weight ratio of $4-1$. The onboard computer is an NVIDIA Orin NX communicating with a Pixracer Pro flight controller running PX4 [40]. All computation is performed onboard the system, and the station only handles visualization and simple initialization commands. This software stack is integrated with ROS2. Our sensing module is an Intel D455 RGB-D Camera. For localization we run either visual-inertial odometry algorithm OpenVINS [41] fused with IMU through an Unscented Kalman Filter (UKF) for indoor flight or GPS module for outdoor flight. Our monocular depth estimation network is DepthAnythingv2[16]. This model is optimized using TensorRT to ensure high speed inference at $80\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$ inference time onboard. To solve our trajectory optimization in eq. (5), we use Interior Point Optimizer (IPOPT) [42] with MA57 [43] as the linear solver. We set a maximum velocity of $2\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$ per axis and similarly a max acceleration of $4\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}^{2}$$ outdoors. This framework uses a SO(3) controller [44] to covert the optimized trajectory to rotation attitude commands for our quadrotor to fly through. During our exploration experiments, we use the Next Best View[39] algorithm to stress our navigation pipeline. Our Next Best View algorithm selects the goals with sample points sampled evenly across the entire map known and unknown regions. We conduct our tests in two areas an indoor flying space with a space of $8\leavevmode\nobreak\ $\mathrm{m}$\times 10\leavevmode\nobreak\ $\mathrm{m}$$, or a larger outdoor flying space $60\leavevmode\nobreak\ $\mathrm{m}$\times 40\leavevmode\nobreak\ $\mathrm{m}$$.
4.2 Outdoor Exploration
Figure 9: Trajectory Path Comparison. Perlin Noise [45] is a gradient noise popular for procedural generation. We took the Perlin Noise Map generator from [11]. Forest Map was borrowed from [21]. The Parking Lot map came from [46]. Hospital was taken from the Amazon Web Services Robomaker.
Figure 10: Trajectory Timing Comparison. Our Method’s Path planning time is under 1 $\mathrm{m}\mathrm{s}$ in all cases therefore is not visible in this plot. Similarly MINCO’s path planning in parking is around 1 $\mathrm{m}\mathrm{s}$ so not visible. SFC refers to the time for safe flight corridor generation. For the map names, we refer to Fig. 9.
Figure 11: Safe flight corridor generation. The height of these layers were manually set to prevent flying over obstacles.
In this set of experiments, we give the quadrotor exploration goal limits to reach $40\%-80\%$ of the declared volume seen in the flight path in Fig. 6. From no map, the quadrotor autonomously explores a region of $60\leavevmode\nobreak\ $\mathrm{m}$\times 40\leavevmode\nobreak\ $\mathrm{m}$$. Overall, we observe that as the planner is tasked with exploring a greater portion of the map, it naturally expands its search path, resulting in a more comprehensive coverage of the environment. Additionally, the vehicle is able to safely navigate in a collision-free manner through the environment in multiple trials showing the repeatability of of the proposed solution. When we combine all the trial flight times, they have a cumulative $20\leavevmode\nobreak\ $\mathrm{m}\mathrm{i}\mathrm{n}$$ of safe collision-free navigation in this environment. It took around $400\leavevmode\nobreak\ $\mathrm{s}$$ for the quadrotor to explore $80\%$ of the environment and about $300\leavevmode\nobreak\ $\mathrm{s}$$ to explore $40\%$. We can analyze the map generated when achieving $80\%$ exploration of the overall space based on its quality in Fig. 7. The map contains all buildings and maintains a similar geometric shape. Additionally, while we observe the presence of some speckle noise in the air, this amount remains minimal. Our tracking error and position following are close to the quadrotor desired value maintaining a tracking error of around $0.2\leavevmode\nobreak\ $\mathrm{m}$$. Although some positional error spikes are observed, they are primarily attributed to environmental factors, as the weather on the testing day included an average wind speed of $15\leavevmode\nobreak\ $\mathrm{m}\mathrm{p}\mathrm{h}$$ with strong $50\leavevmode\nobreak\ $\mathrm{m}\mathrm{p}\mathrm{h}$$ gusts.
4.3 Depth Generation Comparison
Additionally, we demonstrate that our learning-based depth performs far better than the commercial off-the-shelf depth generated by the D455 Stereo Cameras. First, the noise is reduced in the monocular version compared to the stereo. Next, we have an improved effective range in our depth completion. Third, we have a more complete image where structure is preserved. These three properties can be seen based on our experiments in Fig. 8 where we compare the point cloud from our stereo depth camera, Intel D455 which uses a stereo matching algorithm, and our depth completion method. Looking closer at Fig. 8, we notice that the noise is reduced in the left image with the stairs. Depth completion maintains a substantially better structure on the guard rails, stairs, and door. In contrast, the point cloud generated by the D455 depth has a wavelike noise that corrupts the depth’s structure. Additionally, because the walls are self-similar, we see holes in the depth estimates, especially next to the door and bottom window areas that are erroneously considered free space. None of these imperfections exist in the depth completion method proposed in this work. Second, we can see that the range of our depth completion method is improved in the middle image. The initial blue stereo depth only contains the beginning of the L building. On the other hand, because there is enough information on the scale from the initial depth, the depth completion can successfully interpolate the remainder of the building, creating the full structure and $90^{\circ}$ turn. Third, we compare the depth generated in the park. The depth completion point cloud is more complete preserving a greater amount of details. This includes the leaves of the trees, and much more of the park’s buildings are maintained. These features are not contained in the stereo baseline. Additionally, we show in our multimedia supplementary material a full video of the depth completion compared to the stereo depth over a $6\leavevmode\nobreak\ $\mathrm{m}\mathrm{i}\mathrm{n}$$ continuous walk. We notice that not only is the structure of our depth completion good, but also that the scale of the stereo depth image is consistent with depth completion in their common range. In our video, we additionally show that the generated depth is generalizable to both indoor and outdoor scenarios as we take the camera out of the basement. We also timed the average speed of our method as $1.5\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ on the Orin NX showing that not only is the quality improved, but our processing can be done in real-time onboard. Overall, these experiments demonstrate the superiority of our method compared to a standard, product-based stereo vision baseline.
4.4 Trajectory Planning Comparison
We compare our computation time and trajectory path with MINCO [11] which is a well-known global planner used in many similar works for obstacle avoidance works [2, 15] that creates obstacle-free and time-optimized global trajectories. Additionally MINCO [47] has shown to be more computationally efficient and optimal compared to multiple baselines in literature making it an effective benchmark. We compare our methods over $5$ separate maps some taken from popular works [11, 7, 46]. All computation timing and qualitative analysis are performed on an Orin NX to ensure a fair comparison. We record our findings in Fig. 9. Our pipeline takes less time to compute global obstacle free paths especially for complicated maps. This is especially true for the forest and hospital cases where we only take a third of the time of MINCO. Additionally, we can analyze the quality of our paths. While we often take a slightly longer path, our method maintains a similar path length quality to MINCO and contains safety constraints.
We analyze the breakdown of our trajectory planning times in Fig. 10. This includes a path planning, safe flight corridor generation, and finally a trajectory optimization. In our work, we measure the time needed to decompose the map as our safe flight corridor generation, and path planning is the amount of time needed to select among this graph the correct corridors. The results of safe flight corridor generation are shown in Fig. 11. We see the obstacles in gray and the layers of the cuboids in red, green, and blue. The free space is broken down to represent a tight fit around the obstacles. Our work initially assumes that unexplored space is a free space until an obstacle is observed. In MINCO’s, path planning is done using RRT implemented by [48]. Unlike our method, MINCO’s implementation generates SFC along the RRT path alone. In the timing breakdown, our work’s path planning takes under $1\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ making it invisible in Fig. 10 while MINCO path planning takes most of its run time. Additionally, if we compare only the last stage, we notice that our trajectory optimization has a similar completion time. However, we guarantee safety constraints as opposed to MINCO that treats safety as an objective in an unconstrained optimization. Part of the superior speed is based on our trajectory optimization’s objective to minimize the number of optimization variables. MINCO’s path planning objective is based on minimum distance instead. Overall, we show that our method produces faster global trajectories, maintains safety constraints, and is of similar quality to MINCO.
4.5 Indoor Navigation
We show that our method is generalizable to indoor environments. We perform $3$ additional exploration tests presented in Fig. 12. For map $1$ and $2$, the quadrotor explores indoors without GPS and still generates high-quality maps. For map $3$, we demonstrate that our quadrotor can navigate non-convex dead ends using global mapping information. Our drone is able to clear all $3$ scenarios. In the left $2$ images, our exploration path is safe and generates a complete map. Additionally, we see that the generated map reflects the real-world image well containing the mannequin in map $1$ and the barriers in map $2$. In the rightmost image, the robot must navigate behind the barricade. The planner’s height is constrained to prevent it from flying over the obstacle. Furthermore, depth completion has a weakness. Close objects can cause massive distortion in the structure of the image. This is the reason behind the lower quality of map $3$ has with respect to maps $1$ and $2$. Despite these challenging perception and trajectory generation problems, the quadrotor successfully escapes to the other side showing our method’s robustness to sensor corruptions and tight corridors.
5 Conclusion
In this work, we present a complete visual sensing and planning system for autonomous navigation of unknown environments. This method is effective in both indoor and outdoor scenarios and capable of fully autonomous exploration using purely onboard sensors. Our method performs superiorly to other popular off-the-shelf methods in terms of depth sensing, and is computationally faster while maintaining safety guarantees in terms of planning. However, we acknowledge that our method presents a few limitations. When generating our cuboid decomposition, we reprocess the full map each time to create the graph of convex sets. This can become computationally expensive as the map grows in size. In future works, we will convert our algorithm to take incremental updates on the map to reduce the computational burden. Additionally, our depth completion algorithm requires an accurate metric depth to rescale. Even if our method is resilient to a noisy baseline metric depth, it is vulnerable if the baseline depth does not exist. Future works will address these limitations and seek to improve both the depth generation and further reduce the computation time required for trajectory planning.
Figure 12: Indoor Exploration Results: The top images include the generated occupancy map in flight. The bottom images are the color images of our map. The red path is the the path the quadrotor followed. The grayscale represents the height where darker blocks are lower and lighter blocks on top. For improved visualization the floor and ceiling are removed in our maps in post processing.